A call option with an exercise price of $65 will expire in 73 days. No cash...

  1. A call option with an exercise price of $65 will expire in 73 days. No cash payments will be made by the underlying asset over the life of the option. If the underlying asset price is at 70 and the risk-free rate is 5%, the lower bounds for an American Call and European Call should be

    American=5. European=5.63

    American=5.63. European=5

    American=5.63. European=5.63

    American=5. European=5
Answer #1

OPTION C

Lower bound of European Call=Lower bound of American Call=MAX(0,S-X/(1+r)^t)=MAX(0,70-65/1.05^(73/365))=5.6312